one positive number is 6 less than twice a second number, and the product is 140. Find the two numbers

Answer by fractalier(6550)

Call the numbers x and y. Thus we have
xy = 140
and
x = 2y - 6
Now substitute into the first equation and get
(2y - 6)(y) = 140
2y^2 - 6y - 140 = 0
y^2 - 3y - 70 = 0
(y - 10)(y + 7) = 0
y = 10 or y = 7
but it needs to be positive so
y = 10 and then
x = 14